OpenStudy (anonymous):

a geostationary satellite is positioned 35,800 km above earths surface. It takes 24 hours to complete one orbit. The radius of Earth is about 6400 KM. a. what distance does the satellite travel in 1 hour? 3 hours? 2.5 hours? 25 hours? b. after how many hours has the satellite traveled 200,000 KM?

OpenStudy (anonymous):

I dont understand how to find the distance the satellite travels, like I know how to find the radius and circumference but I don't know where to go from there /:

OpenStudy (irishboy123):

"24 hours to complete one orbit": if you can calc the circumference of orbit, then you can use this info to scale down or up and find your answer(s).

OpenStudy (anonymous):

So which circumference would I be taking? 2pi(35,800) or 2pi(6400)? ;-;

OpenStudy (irishboy123):

you need total radius, yes, to get circumference  it is spelt out in the question too.

OpenStudy (wolf1728):

35,800 km above Earth's surface. Earth's radius = 6,400 KM Total orbital radius = 35,800 km plus 6,400 km = 42,200 km orbital circumference = 2 * PI * 42,200 km = 265,150 km. It makes one orbit every 24 hours so every 24 hours it travels 265,150 km. 265,150/24 = 11,048 km per hour I trust that gives you a good start.